OpenStudy (anonymous):

what is the product of (4 sqrrt3-6 sqrrt2)(3 sqrrt3+ 2 sqrrt 2)

OpenStudy (anonymous):

(4rt(3)-6rt(2))(3rt(3)+2rt(2))=36*3+8rt(6)-18rt(6)-12*2

OpenStudy (anonymous):

108-24-10rt(6)

OpenStudy (anonymous):

FOIL: First, Outside, Inside, Last

OpenStudy (anonymous):

Saubhik none of the muiltiple choice answers are this sorry but thanks for trying .. and chaguanas im not sure what you mean by that.. also what i mainly need to know is hwhen i do the square roots times eachother like 3 and 2 square root does that become square root 6

OpenStudy (anonymous):

sorry thats 36-24-10rt(6)=12-10rt(6).

OpenStudy (anonymous):

excuse the silly mistake..

OpenStudy (anonymous):

Thanks that is an answer haha :] im gonna try doing it now

OpenStudy (anonymous):

do it fast and win in life!!

OpenStudy (anonymous):

The question was:\[\text{ what is the product of (4}\sqrt{3}\text{ - 6}\sqrt{2}\text{)(3}\sqrt{3}\text{ + 2}\sqrt{2}\text{) } \]Answer:\[12-10 \sqrt{6} \]